How much would you have to invest now at 6% per year simple interest to have an amount of 3200$ in 2 years
Answer by ewatrrr(24785)

P + I = 3200
1•P + P(.06)(2)= $3200
1•P + .12P= $3200
P(1 + .12) = $3200
P = $3200/1.12 = $2857.14